# Hyperbolic Busemann Neural Networks

### v1

@@ 1 Introduction
Hyperbolic representations have recently delivered strong performance across different applications because the exponential volume growth of negatively curved manifolds enables low-distortion embeddings of tree-like and hierarchical structure [44]. These advantages have been validated in computer vision [22, 32, 18, 54, 23, 4, 29, 3, 52, 39], graph learning [9, 2, 20, 51], multimodal learning [16, 46], recommendation systems [57], astronomy [10], genome sequence learning [31], and natural language processing [44, 21, 45, 25, 28, 55]. These empirical successes highlight the need for principled and efficient hyperbolic neural network components that fully leverage hyperbolic geometry.
@@
Recently, hyperplanes and point-to-hyperplane distances [36] have been adopted to construct hyperbolic MLR in both Poincaré [21, 49, 42] and Lorentz [4] models. Ganea et al. [21, Sec. 3.1] introduced the first Poincaré MLR based on Poincaré hyperplanes, but the formulation suffers from over-parameterization and lacks batch efficiency. Shimizu et al. [49, Sec. 3.1] alleviated these issues through a re-parameterization strategy. Building on these ideas, Bdeir et al. [4, Sec. 4.3] proposed a Lorentz MLR. However, its hyperplane is defined by the ambient Minkowski space, which is model-specific and may distort Lorentzian geometry.
For hyperbolic FC layers, three main formulations exist. Ganea et al. [21, Sec. 3.2] introduced Möbius matrix-vector multiplication through the tangent space on the Poincaré ball. Shimizu et al. [49, Sec. 3.2] further proposed the Poincaré FC layer, defined intrinsically but restricted to the Poincaré model. On the Lorentz model, Chen et al. [11, Sec. 2.2] constructed a Lorentz FC layer by applying linear transformations in the ambient Minkowski space followed by projection onto the Lorentz model. Thus, Möbius and Lorentz FC rely on flat-space (tangent or ambient) approximations that could distort intrinsic geometry, whereas Poincaré FC is intrinsic but model-specific.
On the other hand, the Busemann function and its level sets, horospheres, have emerged as powerful intrinsic tools for hyperbolic learning. They enjoy convenient metric properties [6, Ch. II.8] and admit closed-form expressions on both the Poincaré and Lorentz models [5, Prop. 9]. These operators have supported several hyperbolic algorithms, including SVM [19], PCA [8], Sliced Wasserstein distances [5], and prototype learning [24]. We also note that Nguyen et al. [42, Cor. 4.3] proposed a Poincaré MLR based on the Busemann function. However, its induced point-to-hyperplane distance is pseudo, coincides with the true distance only in Euclidean geometry, remains over-parameterized, and is not batch efficient.
These observations motivate intrinsic and batch-efficient formulations for MLR and FC layers that can operate on both the Poincaré ball and the Lorentz model. To address this need, we propose Busemann Multinomial Logistics Regression (BMLR) and Busemann Fully Connected (BFC), two Busemann-based components for hyperbolic networks. Our contributions are summarized as follows:
@@ 2 Preliminaries
Hyperbolic geometry. An $n$-dimensional hyperbolic space is a Riemannian manifold with constant negative sectional curvature $K<0$. There exist five models of hyperbolic space [7, Sec. 7]. Among them, the Poincaré ball and Lorentz models are the most widely used. The Poincaré ball model [37, p. 62] is $\mathbb{P}^{n}_{K}=\{x\in\mathbb{R}^{n}:\left\|x\right\|^{2}<-1/K\}$ with Riemannian metric $g_{x}(u,v)=(\lambda_{x}^{K})^{2}\left\langle u,v\right\rangle$, where $\lambda_{x}^{K}=\frac{2}{1+K\left\|x\right\|^{2}}$. Here $\left\langle\cdot,\cdot\right\rangle$ and $\left\|\cdot\right\|$ denote the Euclidean inner product and norm. The origin is the zero vector $\mathbf{0}\in\mathbb{P}^{n}_{K}$. The Lorentz model [37, p. 62], also called the hyperboloid model, is $\mathbb{L}^{n}_{K}=\{x\in\mathbb{R}^{n+1}:\left\langle x,x\right\rangle_{\mathcal{L}}=1/K,\ x_{t}>0\}$, where $\left\langle x,y\right\rangle_{\mathcal{L}}=-x_{t}y_{t}+\left\langle x_{s},y_{s}\right\rangle$ is the Lorentzian inner product and $\left\|x\right\|_{\mathcal{L}}=\sqrt{\left\langle x,x\right\rangle_{\mathcal{L}}}$ is the induced Lorentzian norm. Following the convention of special relativity [47, §3.1], we write $x=[x_{t},x_{s}^{\top}]^{\top}$ with temporal component $x_{t}\in\mathbb{R}$ and spatial component $x_{s}\in\mathbb{R}^{n}$. The canonical origin is $\overline{\mathbf{0}}=[\sqrt{1/|K|},0,\ldots,0]^{\top}\in\mathbb{L}^{n}_{K}$. Throughout, we write $\mathcal{H}^{n}_{K}\in\{\mathbb{P}^{n}_{K},\mathbb{L}^{n}_{K}\}$ for a hyperbolic space.
@@
Riemannian & gyrovector operators. The hyperbolic space $\mathcal{H}^{n}_{K}$ provides closed-form expressions for standard Riemannian operators. It also supports a gyrovector structure that extends vector space operations to hyperbolic geometry. On the Poincaré ball $\mathbb{P}^{n}_{K}$, this structure corresponds to the Möbius gyrovector space [53, Ch. 6.14], denoted as $\{\mathbb{P}^{n}_{K},\oplus_{\mathrm{M}},\odot_{\mathrm{M}}\}$. As shown by Chen et al. [13, Props. 24 and 25], the Lorentz model also admits a closed-form gyrovector space, denoted as $\{\mathbb{L}^{n}_{K},\oplus_{\mathbb{L}},\odot_{\mathbb{L}}\}$. Please refer to Sec.˜B.4 for detailed expressions.
@@ 3.1 Formulation
On the Poincaré ball, letting $K\to 0^{-}$ recovers Euclidean geometry [50, App. A.4.2]. For the Lorentz model, as $K\to 0^{-}$, the temporal coordinate diverges while the spatial component approaches $\mathbb{R}^{n}$, making $\mathbb{L}^{n}_{K}$ converge to a Euclidean space. Consistently, the Poincaré and Lorentz Busemann functions and the associated BMLR logits reduce to their Euclidean counterparts, providing a natural generalization of Euclidean MLR.
@@
Poincaré MLR [21, Eq. (25)] $\frac{\lambda_{p_{k}}^{K}\left\|a_{k}\right\|}{\sqrt{-K}}\sinh^{-1}\left(\frac{2\sqrt{-K}\left\langle-p_{k}\oplus_{\mathrm{M}}x,a_{k}\right\rangle}{\left(1+K\left\|-p_{k}\oplus_{\mathrm{M}}x\right\|^{2}\right)\left\|a_{k}\right\|}\right)$, with $p_{k}\in\mathbb{P}^{n}_{K},a_{k}\in T_{p_{k}}\mathbb{P}^{n}_{K}$ $\mathbb{P}^{n}_{K}$ Real $C(2n)$ $C(19n+29)$
Poincaré MLR [49, Eq. (6)] $\frac{2}{\sqrt{-K}}\alpha_{k}\sinh^{-1}\left(\alpha-\beta\right)$, $\alpha=\lambda_{x}^{K}\sqrt{-K}\left\langle x,v_{k}\right\rangle\cosh\left(2\sqrt{-K}b_{k}\right)$, $\beta=\left(\lambda_{x}^{K}-1\right)\sinh\left(2\sqrt{-K}b_{k}\right)$, with $\alpha_{k}>0,v_{k}\in\mathbb{S}^{n-1},b_{k}\in\mathbb{R}$ $\mathbb{P}^{n}_{K}$ Real $C(n+2)$ $C(4n+52)$
Pseudo-Busemann MLR [42, Cor. 4.3] $-d(x,p_{k})\frac{B^{v_{k}}\left(-p_{k}\oplus_{\mathrm{M}}x\right)}{\left\|-p_{k}\oplus_{\mathrm{M}}x\right\|}$, with $p_{k}\in\mathbb{P}^{n}_{K},v_{k}\in\mathbb{S}^{n-1}$ $\mathbb{P}^{n}_{K}$ Pseudo $C(2n)$ $C(19n+34)$
@@ 3.2 Geometric interpretation
As shown by Lebanon and Lafferty [36, Sec. 5], the Euclidean MLR can be reformulated by point-to-hyperplane distances:
@@
Since $B^{v}(x)=-\left\langle v,x\right\rangle$ in Euclidean geometry, Eqs.˜18, 20 and 19 naturally generalize to their Euclidean counterparts. We also acknowledge Fan et al. [19, Eq. (2) and Prop. 3.1], who used horospheres and point-to-horosphere distances to construct a hyperbolic SVM. However, they considered only the unit Poincaré ball with the specific curvature K=1K=-1, which is a special case of our Eqs.˜18 and 20.
@@ 3.3 Comparison
Based on the point-to-hyperplane reformulation in Eq.˜13, recent work extended MLR to the Poincaré [21, 49, 42] and Lorentz [4] models. Ganea et al. [21, Sec. 3.1] introduced the first Poincaré MLR by replacing the Euclidean point-to-hyperplane distance with its hyperbolic counterpart, where the hyperplane is defined by geodesics and the resulting distance is the real point-to-hyperplane distance, obtained as an infimum over the hyperplane. However, the formulation is not batch efficient (see App.˜C). It also requires per-class parameters $a_{k}\in T_{p_{k}}\mathbb{P}^{n}_{K}$ and $p_{k}\in\mathbb{P}^{n}_{K}$, which leads to over-parameterization. Shimizu et al. [49, Sec. 3.1] alleviated such issues via re-parameterization. Bdeir et al. [4, Sec. 4.3] further developed a Lorentz MLR, but its hyperplanes are defined by the ambient Minkowski space, which is tailored to the Lorentz model and does not fully respect the intrinsic hyperbolic geometry. Moreover, Nguyen et al. [42, Cor. 4.3] proposed a Poincaré MLR based on the Busemann function. We refer to it as Pseudo-Busemann MLR, as the induced point-to-hyperplane distance is pseudo, coinciding with the real point-to-hyperplane distance only in Euclidean geometry. It also suffers from over-parameterization and is not batch efficient.
As summarized in Tab.˜2111Relative to [42, Def. 4.2, Cor. 4.3, and App. B.1.2], the Pseudo-Busemann MLR written here includes an additional sign -; this is intentional and matches their official implementation., BMLR unifies advantages that prior hyperbolic MLRs offer only partially. In particular, BMLR respects the real point-to-horosphere distance, uses compact parameters without an additional manifold-valued point, attains the lowest FLOPs on $\mathbb{L}^{n}_{K}$ and a competitive cost on $\mathbb{P}^{n}_{K}$, and supports batch-efficient computation. On $\mathbb{L}^{n}_{K}$, its FLOPs are even close to those of the Euclidean MLR.
@@ 4.1 Formulation
Euclidean FC layers. An FC affine transformation $\mathcal{F}:\mathbb{R}^{n}\ni x\mapsto y=Ax+b\in\mathbb{R}^{m}$ can be expressed element-wise as $y_{k}=\left\langle a_{k},x\right\rangle+b_{k}$ with $a_{k}\in\mathbb{R}^{n}$ and $b_{k}\in\mathbb{R}$. As shown by Shimizu et al. [49, Sec. 3.2], the LHS $y_{k}$ can be interpreted as the signed distance from $y$ to the hyperplane passing through the origin and orthogonal to the $k$-th axis of the output space. Hence, the FC layer can be expressed as
@@
Lifting to hyperbolic space. To extend Eq.˜21 into hyperbolic space, the RHS can be readily replaced by Eq.˜8, as it generalizes the term $\left\langle a_{k},x\right\rangle+b_{k}$. For the LHS, a natural idea is to use the signed point-to-horosphere distance. However, as detailed in App.˜D, this may fail to admit a solution for $y$. We therefore follow the point-to-hyperplane distance in [21, Thm. 5] for the Poincaré model and the one in [4, Eq. (44)] for the Lorentz model. Given $x\in\mathcal{H}^{n}_{K}$, the hyperbolic BFC layer $\mathcal{F}:\mathcal{H}^{n}_{K}\ni x\mapsto y\in\mathcal{H}^{m}_{K}$ is given by solving $y$ via the following $m$ equations:
@@
The proof, together with $H_{e_{k},e}$ and $\bar{\operatorname{d}}\left(y,H_{e_{k},e}\right)$, is provided in Sec.˜F.3, which is inspired by Shimizu et al. [49, App. D.3]. ∎
@@
Möbius [21, Eq. 27] $\frac{1}{\sqrt{-K}}\tanh\left(\frac{\left\|Wx\right\|}{\left\|x\right\|}\tanh^{-1}\left(\sqrt{-K}\left\|x\right\|\right)\right)\frac{Wx}{\left\|Wx\right\|}$ $\mathbb{P}^{n}_{K}$ Tangent $W\in\mathbb{R}^{m\times n}$ mnmn $2nm+2n$ +2m+24+2m+24
Poincaré FC [49, Eq. (7)] $y=\frac{\omega}{1+\sqrt{1-K\left\|\omega\right\|^{2}}},\ \omega_{k}=\frac{\sinh\left(\sqrt{-K}u_{k}(x)\right)}{\sqrt{-K}}$, with $u_{k}(x)$ in Tab.˜2 $\mathbb{P}^{n}_{K}$ Poincaré geometry $\alpha_{k}>0,\ v_{k}\in\mathbb{S}^{n-1}$, $b_{k}\in\mathbb{R}$, for $k=1,\ldots,m$ $m(n+2)$ $4nm+71m+4$

### v2

@@ 1 Introduction
Hyperbolic representations have recently delivered strong performance across different applications because the exponential volume growth of negatively curved manifolds enables low-distortion embeddings of tree-like and hierarchical structure [46]. These advantages have been validated in computer vision [24, 34, 20, 56, 25, 4, 31, 3, 54, 42, 57], graph learning [9, 2, 22, 53], multimodal learning [18, 48], recommendation systems [60], astronomy [10], genome sequence learning [33], natural language processing [46, 23, 47, 27, 30, 58], and brain signal decoding [40]. These empirical successes highlight the need for principled and efficient hyperbolic neural network components that fully leverage hyperbolic geometry.
@@
Recently, hyperplanes and point-to-hyperplane distances [38] have been adopted to construct hyperbolic MLR in both Poincaré [23, 51, 45] and Lorentz [4] models. Ganea et al. [23, Sec. 3.1] introduced the first Poincaré MLR based on Poincaré hyperplanes, but the formulation suffers from over-parameterization and lacks batch efficiency. Shimizu et al. [51, Sec. 3.1] alleviated these issues through a re-parameterization strategy. Building on these ideas, Bdeir et al. [4, Sec. 4.3] proposed a Lorentz MLR. However, its hyperplane is defined by the ambient Minkowski space, which is model-specific and may distort Lorentzian geometry.
For hyperbolic FC layers, three main formulations exist. Ganea et al. [23, Sec. 3.2] introduced Möbius matrix-vector multiplication through the tangent space on the Poincaré ball. Shimizu et al. [51, Sec. 3.2] further proposed the Poincaré FC layer, defined intrinsically but restricted to the Poincaré model. On the Lorentz model, Chen et al. [11, Sec. 2.2] constructed a Lorentz FC layer by applying linear transformations in the ambient Minkowski space followed by projection onto the Lorentz model. Thus, Möbius and Lorentz FC rely on flat-space (tangent or ambient) approximations that could distort intrinsic geometry, whereas Poincaré FC is intrinsic but model-specific.
On the other hand, the Busemann function and its level sets, horospheres, have emerged as powerful intrinsic tools for hyperbolic learning. They enjoy convenient metric properties [6, Ch. II.8] and admit closed-form expressions on both the Poincaré and Lorentz models [5, Prop. 9]. These operators have supported several hyperbolic algorithms, including SVM [21], PCA [8], Sliced Wasserstein distances [5], and prototype learning [26]. We also note that Nguyen et al. [45, Cor. 4.3] proposed a Poincaré MLR based on the Busemann function. However, its induced point-to-hyperplane distance is pseudo, coincides with the true distance only in Euclidean geometry, remains over-parameterized, and is not batch efficient.
These observations motivate intrinsic and batch-efficient formulations for MLR and FC layers that can operate on both the Poincaré ball and the Lorentz model. To address this need, we propose Busemann Multinomial Logistics Regression (BMLR) and Busemann Fully Connected (BFC) layers, two Busemann-based components for hyperbolic networks. Our contributions are summarized as follows:
@@ 2 Preliminaries
Hyperbolic geometry. An $n$-dimensional hyperbolic space is a Riemannian manifold with constant negative sectional curvature $K<0$. There exist five models of hyperbolic space [7, Sec. 7]. Among them, the Poincaré ball and Lorentz models are the most widely used. The Poincaré ball model [39, p. 62] is $\mathbb{P}^{n}_{K}=\{x\in\mathbb{R}^{n}:\left\|x\right\|^{2}<-1/K\}$ with Riemannian metric $g_{x}(u,v)=(\lambda_{x}^{K})^{2}\left\langle u,v\right\rangle$, where $\lambda_{x}^{K}=\frac{2}{1+K\left\|x\right\|^{2}}$. Here $\left\langle\cdot,\cdot\right\rangle$ and $\left\|\cdot\right\|$ denote the Euclidean inner product and norm. The origin is the zero vector $\mathbf{0}\in\mathbb{P}^{n}_{K}$. The Lorentz model [39, p. 62], also called the hyperboloid model, is $\mathbb{L}^{n}_{K}=\{x\in\mathbb{R}^{n+1}:\left\langle x,x\right\rangle_{\mathcal{L}}=1/K,\ x_{t}>0\}$, where $\left\langle x,y\right\rangle_{\mathcal{L}}=-x_{t}y_{t}+\left\langle x_{s},y_{s}\right\rangle$ is the Lorentzian inner product and $\left\|x\right\|_{\mathcal{L}}=\sqrt{\left\langle x,x\right\rangle_{\mathcal{L}}}$ is the induced Lorentzian norm. Following the convention of special relativity [49, §3.1], we write $x=[x_{t},x_{s}^{\top}]^{\top}$ with temporal component $x_{t}\in\mathbb{R}$ and spatial component $x_{s}\in\mathbb{R}^{n}$. The canonical origin is $\overline{\mathbf{0}}=[\sqrt{1/|K|},0,\ldots,0]^{\top}\in\mathbb{L}^{n}_{K}$. Throughout, we write $\mathcal{H}^{n}_{K}\in\{\mathbb{P}^{n}_{K},\mathbb{L}^{n}_{K}\}$ for a hyperbolic space.
@@
Riemannian & gyrovector operators. The hyperbolic space $\mathcal{H}^{n}_{K}$ provides closed-form expressions for standard Riemannian operators. It also supports a gyrovector structure that extends vector space operations to hyperbolic geometry. On the Poincaré ball $\mathbb{P}^{n}_{K}$, this structure corresponds to the Möbius gyrovector space [55, Ch. 6.14], denoted as $\{\mathbb{P}^{n}_{K},\oplus_{\mathrm{M}},\odot_{\mathrm{M}}\}$. As shown by Chen et al. [15, Props. 24 and 25], the Lorentz model also admits a closed-form gyrovector space, denoted as $\{\mathbb{L}^{n}_{K},\oplus_{\mathbb{L}},\odot_{\mathbb{L}}\}$. Please refer to Sec.˜B.4 for detailed expressions.
@@ 3.1 Formulation
On the Poincaré ball, letting $K\to 0^{-}$ recovers Euclidean geometry [52, App. A.4.2]. For the Lorentz model, as $K\to 0^{-}$, the temporal coordinate diverges while the spatial component approaches $\mathbb{R}^{n}$, making $\mathbb{L}^{n}_{K}$ converge to a Euclidean space. Consistently, the Poincaré and Lorentz Busemann functions and the associated BMLR logits reduce to their Euclidean counterparts, providing a natural generalization of Euclidean MLR.
@@
Poincaré MLR [23, Eq. (25)] $\frac{\lambda_{p_{k}}^{K}\left\|a_{k}\right\|}{\sqrt{-K}}\sinh^{-1}\left(\frac{2\sqrt{-K}\left\langle-p_{k}\oplus_{\mathrm{M}}x,a_{k}\right\rangle}{\left(1+K\left\|-p_{k}\oplus_{\mathrm{M}}x\right\|^{2}\right)\left\|a_{k}\right\|}\right)$, with $p_{k}\in\mathbb{P}^{n}_{K},a_{k}\in T_{p_{k}}\mathbb{P}^{n}_{K}$ $\mathbb{P}^{n}_{K}$ Real $C(2n)$ $C(19n+29)$
Poincaré MLR [51, Eq. (6)] $\frac{2}{\sqrt{-K}}\alpha_{k}\sinh^{-1}\left(\alpha-\beta\right)$, $\alpha=\lambda_{x}^{K}\sqrt{-K}\left\langle x,v_{k}\right\rangle\cosh\left(2\sqrt{-K}b_{k}\right)$, $\beta=\left(\lambda_{x}^{K}-1\right)\sinh\left(2\sqrt{-K}b_{k}\right)$, with $\alpha_{k}>0,v_{k}\in\mathbb{S}^{n-1},b_{k}\in\mathbb{R}$ $\mathbb{P}^{n}_{K}$ Real $C(n+2)$ $C(4n+52)$
Pseudo-Busemann MLR [45, Cor. 4.3] $-d(x,p_{k})\frac{B^{v_{k}}\left(-p_{k}\oplus_{\mathrm{M}}x\right)}{\left\|-p_{k}\oplus_{\mathrm{M}}x\right\|}$, with $p_{k}\in\mathbb{P}^{n}_{K},v_{k}\in\mathbb{S}^{n-1}$ $\mathbb{P}^{n}_{K}$ Pseudo $C(2n)$ $C(19n+34)$
@@ 3.2 Geometric interpretation
As shown by Lebanon and Lafferty [38, Sec. 5], the Euclidean MLR can be reformulated by point-to-hyperplane distances:
@@
Since $B^{v}(x)=-\left\langle v,x\right\rangle$ in Euclidean geometry, Eqs.˜18, 20 and 19 naturally generalize to their Euclidean counterparts. We also acknowledge Fan et al. [21, Eq. (2) and Prop. 3.1], who used horospheres and point-to-horosphere distances to construct a hyperbolic SVM. However, they considered only the unit Poincaré ball with the specific curvature K=1K=-1, which is a special case of our Eqs.˜18 and 20.
@@ 3.3 Comparison
Based on the point-to-hyperplane reformulation in Eq.˜13, recent work extended MLR to the Poincaré [23, 51, 45] and Lorentz [4] models. Ganea et al. [23, Sec. 3.1] introduced the first Poincaré MLR by replacing the Euclidean point-to-hyperplane distance with its hyperbolic counterpart, where the hyperplane is defined by geodesics and the resulting distance is the real point-to-hyperplane distance, obtained as an infimum over the hyperplane. However, the formulation is not batch efficient (see App.˜C). It also requires per-class parameters $a_{k}\in T_{p_{k}}\mathbb{P}^{n}_{K}$ and $p_{k}\in\mathbb{P}^{n}_{K}$, which leads to over-parameterization. Shimizu et al. [51, Sec. 3.1] alleviated such issues via re-parameterization. Bdeir et al. [4, Sec. 4.3] further developed a Lorentz MLR, but its hyperplanes are defined by the ambient Minkowski space, which is tailored to the Lorentz model and does not fully respect the intrinsic hyperbolic geometry. Moreover, Nguyen et al. [45, Cor. 4.3] proposed a Poincaré MLR based on the Busemann function. We refer to it as Pseudo-Busemann MLR, as the induced point-to-hyperplane distance is pseudo, coinciding with the real point-to-hyperplane distance only in Euclidean geometry. It also suffers from over-parameterization and is not batch efficient.
As summarized in Tab.˜2111Relative to [45, Def. 4.2, Cor. 4.3, and App. B.1.2], the Pseudo-Busemann MLR written here includes an additional sign -; this is intentional and matches their official implementation., BMLR unifies advantages that prior hyperbolic MLRs offer only partially. In particular, BMLR respects the real point-to-horosphere distance, uses compact parameters without an additional manifold-valued point, attains the lowest FLOPs on $\mathbb{L}^{n}_{K}$ and a competitive cost on $\mathbb{P}^{n}_{K}$, and supports batch-efficient computation. On $\mathbb{L}^{n}_{K}$, its FLOPs are even close to those of the Euclidean MLR.
@@ 4.1 Formulation
Euclidean FC layers. An FC affine transformation $\mathcal{F}:\mathbb{R}^{n}\ni x\mapsto y=Ax+b\in\mathbb{R}^{m}$ can be expressed element-wise as $y_{k}=\left\langle a_{k},x\right\rangle+b_{k}$ with $a_{k}\in\mathbb{R}^{n}$ and $b_{k}\in\mathbb{R}$. As shown by Shimizu et al. [51, Sec. 3.2] and Chen et al. [14, Sec. 3.1], the LHS $y_{k}$ can be interpreted as the signed distance from $y$ to the hyperplane passing through the origin and orthogonal to the $k$-th axis of the output space. Hence, the FC layer can be expressed as
@@
Lifting to hyperbolic space. To extend Eq.˜21 into hyperbolic space, the RHS can be readily replaced by Eq.˜8, as it generalizes the term $\left\langle a_{k},x\right\rangle+b_{k}$. For the LHS, a natural idea is to use the signed point-to-horosphere distance. However, as detailed in App.˜D, this may fail to admit a solution for $y$. We therefore follow the point-to-hyperplane distance in [23, Thm. 5] for the Poincaré model and the one in [4, Eq. (44)] for the Lorentz model. Given $x\in\mathcal{H}^{n}_{K}$, the hyperbolic BFC layer $\mathcal{F}:\mathcal{H}^{n}_{K}\ni x\mapsto y\in\mathcal{H}^{m}_{K}$ is given by solving $y$ via the following $m$ equations:
@@
The proof, together with $H_{e_{k},e}$ and $\bar{\operatorname{d}}\left(y,H_{e_{k},e}\right)$, is provided in Sec.˜F.3, which is inspired by Shimizu et al. [51, App. D.3]. ∎
@@
Möbius [23, Eq. 27] $\frac{1}{\sqrt{-K}}\tanh\left(\frac{\left\|Wx\right\|}{\left\|x\right\|}\tanh^{-1}\left(\sqrt{-K}\left\|x\right\|\right)\right)\frac{Wx}{\left\|Wx\right\|}$ $\mathbb{P}^{n}_{K}$ Tangent $W\in\mathbb{R}^{m\times n}$ mnmn $2nm+2n$ +2m+24+2m+24
Poincaré FC [51, Eq. (7)] $y=\frac{\omega}{1+\sqrt{1-K\left\|\omega\right\|^{2}}},\ \omega_{k}=\frac{\sinh\left(\sqrt{-K}u_{k}(x)\right)}{\sqrt{-K}}$, with $u_{k}(x)$ in Tab.˜2 $\mathbb{P}^{n}_{K}$ Poincaré geometry $\alpha_{k}>0,\ v_{k}\in\mathbb{S}^{n-1}$, $b_{k}\in\mathbb{R}$, for $k=1,\ldots,m$ $m(n+2)$ $4nm+71m+4$